OpenStudy (anonymous):

translate into an equation Seven times the sum of a number and 8 equals 9 use the variable x for the unknown number

OpenStudy (anonymous):

7(X+8) = 9 7X + 56 = 9 7X = -47 X = -47/7

OpenStudy (anonymous):

ty